use a unit multiplyer to convert 11 inches to centimeters.(1in. = 2..54cm.) {the answers are..........2.794cm., 27.94cm., 4.33cm., or none of these}
Answer by askmemath(368)

If 1 inch = 2.54cm
11inches = 11x2.54 = 27.94cms
Here's a trick for multiplying with 11.Starting from the right side,write down the right most digit in this case 4. Then,add the rightmost digit and the digit to the left of it 4+5 = 9(If its greater than 9,write the left digit and carry forward).Now we add 2+5 = 7.And finally we write the left most digit as it is in this case 2. Since there are two digits after the decimal place,we place the decimal point 2 places from the right.